OpenStudy (anonymous):

What is the solution to the equations? 4x+2y=6 2x+y=3

OpenStudy (omgthatgirl):

multiple choice?

OpenStudy (dtan5457):

your looking for a solution for x and y i'd do this by elimination 4x+2y=6 2x+y=3<<<multiply by -2 -4x-2y=-6 4x+2y=6 add the, what do you get?

OpenStudy (anonymous):

I just don't get this stuff at all. I believe the 2 equations end up having the same answer though

OpenStudy (dtan5457):

exactly! infinite solutions

OpenStudy (anonymous):

Y= 3-2x and x = 3/2-1/2y

OpenStudy (dtan5457):

if they are the same line, when graphed, they will infinite intersections. therefore infinite solutions

OpenStudy (anonymous):

ok dtan. thanks for your math insight, you've been helpful.